What is the lowest acceptable probability number beyond which it would considered impossible?
Answer 1: Borel's Law tells us that anything with a probability less than 1 in 1050 is "mathematically impossible." Leslie Orgel calculated that any event that has one chance in just 10^50 (1 followed by 50 zeros) is dismissed by mathematicians as never happening. Mathematician William Dembski's calculations can be summed up with the following simplification of the primary principle: "If you roll a 6 sided die, one time, what are the chances that you will roll a 7 or higher?" The correct answer is that it's impossible. Thus, the resulting number can not exceed the afforded chances.
Answer 2: An impossible event is if there is a 0% chance of it occurring. If you were to flip a coin until you got 'cheese' as a result, then there is a 0% chance, it is not possible.
Above 0, nothing is impossible, they just have greater and greater degrees of improbability. Tossing 25 coins and getting 10 heads in a row for example is 1/33554432 or 0.00000298% which sounds impossible. But if you got 33,554,432 people to simultaneously flip 25 coins each, in theory one person should get 25 heads.
What is the probability of rolling a 3 or a 4 with a six sided die?
The probability of rolling a 3 or 4 with a 6 sided die is 2 over 6. Also known as one/third.
What is the probability of getting a seven of spades?
I t is necessary to assume that
If these assumptions are realised - and there is nothing in the question to suggest that they are - then the probability is 1/52.
What is the probability of a spinner not landing on an even number 1-6?
If I have understood the question correctly, the answer is 1/2.
The risk.
Is the number of all possible outcomes of an experiment. The number depends on the experiment.
What is the probability that Alex makes both of his shots?
The answer depends on the probability that Alex makes a single shot.
How many different ways can you arrange 5 letters in a row?
5 choices for the first, 4 for the second .... so is 5x4x3x2x1 = 120 ways. If 2 of the letters are the same then for any arrangement, switching those 2 doesn't make any difference, so divide the answer by 2. If 3 are the same then there are 3x2x1 ways of shuffling those, which at first were all different arrangements but now are not, so divide by 3x2x1 ie 6.
Is homogeneous uniform distribution?
Your question cannot be answered as stated. Please explain further.
Given the way you have worded the question I take it to mean, what is the probability of drawing at least one spade?
We can do this most easily by asking first, what is the probability of drawing no spades on each of the 80 times. This is 39/52. The probability of doing this 80 times is (39/52)80.
Then the probability of not doing this is 1 - (39/52)80, which is quite close to one. The probability of drawing at least one spade is almost one.
What is the probability of an impossible event if zero is empirical?
The probabilit of an impossible event is 0. Empiricism has nothing whatsoever to do with it.
50%. 1 of the two male offspring will definately be color blind. Do a punnett square with the father having normal vision and the mother being a carrier.
What is the probability that a standard normal variable will be between the values of -2 and 1?
0.636 approx.
What is the probability of drawing a diamond and a black seven in a deck of 52 cards?
# of diamond card = 13
# of black 7 card = 2
P(diamond and black 7) = 13/52 * 2/51 = 1/4 * 2/51
Answer = 1/102
0.8.
The probability that one die will not be a 3 is 5/6
The probability that all will not be 3 is ((5/6)4)
If you roll 4 dice, there are 5 possible outcomes (for this problem).
There will be NO 3's, there will be one 3, there will be two 3's. there will be three 3's, or, there will be four 3's. Add all possible probabilities = 1. At least one meanse either 1,2,3, or 4 3's but NOT one.
So the prob of at least one = total prob - prob of NO
1- (5/6)4 = 1 - 625/1296
= 671/1296 <-- answer
How many possible ways to answer three true or false questions?
2 possibilities: true or false
3 questions
2^3
2*2*2 = 8 ways
What is the probability that 4 and 7 randomly chosen numbers in a row are all divisible by 7?
4 numbers: 1/74 = 0.000 061 035
7 numbers: 1/77 = 0.000 001 214